OpenStudy (anonymous):

PLEASE HELP!!!!! Add and identify any x-values for which the following expression is undefined: (3x-2)/(x+2) + (2x)/(4x-1) Thanks! :)

OpenStudy (anonymous):

Ok, so the expression is undefined if you divide by 0.

OpenStudy (anonymous):

So either (x+2) = 0 or (4x-1) = 0

OpenStudy (anonymous):

so x = -2, or x = 1/4

OpenStudy (anonymous):

Okay, thanks! That part makes sense!

OpenStudy (anonymous):

So the values of x for which the expression is undefined are -2, 0.25

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

But!

OpenStudy (anonymous):

they want us to add first, so you have to find a common denominator and add the fractions

OpenStudy (anonymous):

(4x-1)(x+2)

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

So what is the new fraction?

OpenStudy (anonymous):

(12x^2-8x-3x+2+2x^2+4x)/(4x-1)(x+2)

OpenStudy (anonymous):

Simplified: (14x^2-7x+2)/(4x-1)(x+2)

OpenStudy (anonymous):

Good, and now, this expression is undefined when the denominator is 0

OpenStudy (anonymous):

Can the numerator be factored any more and then be canceled with one of the factors from the denominator?

OpenStudy (anonymous):

In this case, no. But even if you could, 0/0 is undefined (IT IS NOT EQUAL TO 1!), so you could not cancel them anyway

OpenStudy (anonymous):

Okay, great! Thanks so much - just the answer I was looking for! :)

OpenStudy (anonymous):

:-) Good luck